### v1

# Towards Stable and Faithful Inpainting

Abstract
Recent $*{}^{*}$: Equal contribution. ${}^{\dagger}$: Corresponding author. in inpainting increasingly relies on generative models, leveraging their strong generation capabilities for addressing ill-conditioned problems. However, this enhanced generation often introduces instability, leading to arbitrary object generation within masked regions. This paper proposes a balanced solution, emphasizing the importance of unmasked regions in guiding inpainting while preserving generative capacity. Our approach, Aligned Stable Inpainting with UnKnown Areas Prior (ASUKA), employs a reconstruction-based masked auto-encoder (MAE) as a stable prior. Aligned with the robust Stable Diffusion inpainting model (SD), ASUKA significantly improves inpainting stability. ASUKA further aligns masked and unmasked regions through an inpainting-specialized decoder, ensuring more faithful inpainting. To validate effectiveness across domains and masking scenarios, we evaluate on MISATO, a collection of several existing dataset. Results confirm ASUKA’s efficacy in both stability and fidelity compared to SD and other inpainting algorithms.
Figure 1: Comparison between ASUKA and reconstruction-based (MAE [20]) as well as generative-based (MAE-FAR [8], MAT [26], Co-Mod [61], LaMa [45], Stable Diffusion Inpainting Model [39]) inpainting algorithms. MAE offers stable masked region estimates, yet falls short in texture detail. GAN-based inpainting struggles with low fidelity, such as omitted white horizontal lines. SD is powerful but unstable, often introducing random elements and suffers from mask-unmask color inconsistency. ASUKA ensures consistency of masked-unmasked areas during diffusion and decoding processes, achieving stable and faithful inpainting. This image is from COCO [30].
1 Introduction
Image inpainting [5] generates content to fill in masked areas of images while maintaining consistency with the unmasked areas. Traditional inpainting algorithms [5, 14, 19, 25, 40] primarily rely on information from unmasked regions to reconstruct masked regions, resulting stable estimations but lacks fidelity. Moreover, the conventional reconstruction loss is prone to generating blurred synthesis due to the inherent ill-posed inpainting issues [36].
To facilitate inpainting with complex mask structures, recent approaches [35, 28, 7, 45, 46, 61] widely utilize the generative models [18, 22, 16], which introduce randomness into the inpainting process, like learning connections between image and noise distributions [18, 22] and utilizing Gibbs sampling [55, 11]. These induced stochasticity alleviates the ill-posed issues, yielding more flexible and faithful generation results [61, 62, 26].
However, every benefit comes with a cost. Current approaches mainly use random masking strategy to train inpainting algorithms, which usually leads to masking foreground objects. When trained with these masks, the model learns to generate foreground objects without information of the object, but sorely based on unrelated background. This introduces a bias in generative models, causing them to hallucinate unreasonable objects in the masked areas. This contradicts the goal of achieving stable inpainting.
In this paper, we propose to combine the strengths of both regression-based reconstruction and distribution-aligned generation to achieve stable and faithful inpainting simultaneously. Particularly, we adopt the Masked Auto-Encoder (MAE) [20] and Stable Diffusion inpainting model (SD) [39], respectively. MAE is trained simply to estimate masked pixels through the L2 reconstruction loss, resulting in stable and average results, albeit with some blurring. On the other hand, SD has impressive generation capacity, but may generate implausible content. In this paper, we demonstrate that the exceptionally faithful feature priors from MAEs can substantially enhance the inpainting stability of SD without sacrificing the performance as shown in Figure 1 with a frozen SD.
Particularly, we manipulate the generation and decoding processes of SD to achieve stable and faithful inpainting with the information from unmasked region.
i) In the generation process, we use MAE to estimate masked region sorely from unmasked region. Specifically, we remove the text condition and turn SD into a MAE prior guided image inpainting model. To align SD with MAE while keeping SD frozen, we propose an alignment module, which is trained on MAE-prior-to-image task with standard diffusion objective and achieves dimension and distribution alignment to cross-attention layers of SD.
ii) In the decoding process, we ensure consistency between the masked and unmasked regions. The use of Variational Auto-Encoder trained with KL divergence loss (KL-VAE) in SD inevitably leads to degradation, causing the color inconsistencies in some inpainting cases. Zhu et al. [64] observed similar issue and train a decoder conditioned on the unmasked regions. However, their decoder only preserves the unmasked region and still suffers from the inconsistency between the masked and unmasked regions. To tackle this challenge, we re-formulate the decoder of KL-VAE as a local harmonization model conditioned on unmasked image pixels. Such a specialized inpainting decoder significantly improves the perceptual consistency of inpainted images.
Combined together, our proposed system, dubbed Aligned Stable inpainting with Unknown Areas prior (ASUKA), is illustrated in Figure 2. To evaluate the effectiveness of inpainting algorithms across various scenarios and mask shapes, we utilize an evaluation dataset named MISATO, which selects representative testing images from Matterport3D [10], Flickr-Landscape [29], MegaDept[27], and COCO 2014 [30]. This dataset thus covers four distinct domains—landscape, indoor, building, and background—making it diverse enough to serve as a benchmark for evaluation. Through experiments conducted on MISATO, especially with large irregular masks, we validate the efficacy of ASUKA.
Contributions. Our contributions can be summarized as:
  • We emphasize stability in inpainting tasks as the generative models suffers from the instability problem;
  • We balance fidelity and stability through aligning reconstruction-based Masked Auto-Encoder (MAE) and generation-based Stable Diffusion (SD).
  • We ensure mask-unmask consistency in the decoding process of SD by re-formulating decoding as a local harmonization task conditioned on unmasked image pixels and training a inpainting-specialized decoder.
2 Related Works
Image inpainting is a longstanding challenge to infill the missing regions with consistent pixels. Classical approaches, relying on patchmatch [15, 4, 58] and partial differential equation [5, 9, 6] are constrained by low-level image features and often fail to produce reasonable results for large holes. Benefits from the adversarial training [18], learning-based GANs achieve prominent success in image inpainting [36, 54, 61, 26, 8]. In particular, GAN-based inpainting methods have evolved with adaptive convolutions [31, 54, 57], attention mechanism [53, 52, 56, 24], and high-resolution generalization in frequency learning [45, 51, 13]. Methods like Co-Mod [61] prioritize adversarial learning and plausible generation to address the challenging ill-posed inpainting issue [26, 62], yielding robust and high-quality generations, even for large masks. However, they suffer from unstable results and uncontrollable hallucinations conditioned on different random latent codes. On the other hand, methods emphasizing higher reconstruction penalties [35, 45, 8] achieve more stable generations, but struggle with large masks, often resulting in blur artifacts. Recent diffusion models [41, 39] have rapidly developed and demonstrated impressive inpainting results. However, these models share similarities with GAN-based methods, learning to align distributions rather than achieving faithful pixel space reconstruction. As a consequence, these diffusion-based approaches face challenges in maintaining faithful inpainting results.
Adapting Stable Diffusion. Currently, SD [39] stands out as one of the most widely used diffusion models. SD incorporates a KL-VAE to learn a latent space preserving image semantics at a lower resolution, along with a U-Net for diffusion within this latent space. Various tuning and aligning methods have been developed to introduce additional conditions to SD. This includes image-inversion for text-guided image-to-image translation [34], textual-inversion for personalization [17], LoRA fine-tuning [23], and controlnet [59] to add other types of conditions. In this paper, our objective is to enhance stability while preserving the generation capacity of the stable diffusion model. Consequently, we refrain from fine-tuning the U-Net backbone of SD. Conversely, the text-condition is not utilized in inpainting tasks. Textual-inversion [17] introduces learnable tokens in Clip text encoders to improve subject description. ObjectStitch [43] employs the CLIP visual feature to represent the subject. In our approach, we advocate removing the text-condition and instead guiding SD using the Masked Auto-Encoder (MAE) prior for masked region.
Information loss in SD latent space. Although claimed only eliminates imperceptible details, the KL-VAE used by SD causes distortion in the reconstruction of images. OpenAI [2] proposes a larger decoder to improve the decoding quality of SD’s latent. Luo et al. [33] propose a frequency-augmented decoder to address the super-resolution case. Zhu et al. [64] propose to preserve unmasked regions during decoding. In this paper, inspired by these approaches, we take a step further to ensure the low-frequency semantic consistency between mask and unmask regions.
Masked Image-Modeling [3] (MIM) is an active research area in self-supervised learning. Typical MIM methods [3, 20, 50, 12] split images into visible and masked patches, learning to estimate masked patches from visible patches. Training targets for visible patches encompass pixel values [20], HOG features [48], and high-level semantic features [49]. While the primary objective of MIM is representation learning, its potential effectiveness in image generation is also noteworthy. Cao et al. [8] adopts MAE features and attention scores to assist the convolutional inpainting model better in handling long-distance dependencies. In contrast, this paper proposes facilitating MAE as the inpainting prior for SD and enhancing the stability of the inpainting process by incorporating MAE prior guidance into the cross-attention modules of SD.
3 Methodology
Problem setup. Inpainting takes as inputs a masked image to complete with a mask to indicate the missing region. The target of inpainting is to fill the missing region based on the information of unmasked regions to generate high-fidelity images. In this paper, we focus on the standard inpainting task without utilizing other conditions.
Figure 2: ASUKA adopts the MAE to provide a low-resolution stable prior for frozen SD to maintain the generation capacity while increasing stability. A better decoder is utilized to achieve mask-unmask semantic consistency when decoding SD results from latent to image for better fidelity.
Overview. The framework of the proposed Aligned Stable inpainting with Unknown Areas prior (ASUKA) is illustrated in Figure 2. ASUKA adopts the pre-trained stable diffusion inpainting model (SD) [39]. Our target is to provide more stable and faithful inpainting results while fully exploiting the generation capacity of SD. Our alignment includes i) a stable alignment to align stable Masked Auto-Encoder (MAE) prior for masked region with SD and ii) a faithful alignment to align ground-truth unmasked region with generated masked region during decoding. To this end, we freeze the U-Net of SD, while replacing the text-condition part with our proposed MAE prior to increase stability. To align the MAE prior to SD, we introduce a alignment module, which is trained via the training objective of SD. Additionally, to align masked and unmasked regions during decoding and resolve the problem of information loss from the KL-VAE decoder which causes mask-unmask inconsistency to SD, we train an inpainting-specialized decoder to better decode the latent back to the image space for seamless fidelity. Combined together, ASUKA achieves more stable and faithful inpainting results.
3.1 Masked Auto-Encoder Prior
Figure 3: Use MAE prior for image-to-image translation (start from 80% noise rate) via SD achieves poor inpainting results.
MAE as stable prior. While recent generative models rely on stochastic noise to provide more diverse generation results, it leads to the generation of random objects unexpectedly. On the other hand, MAE is known to provide a stable estimation of masked regions based purely on the unmasked regions. As MAE is trained on the L2 reconstruction loss, we can regard the estimation of MAE as a mean estimation, which can be utilized to provide a stable prior for generative models to not generate new concepts. However, MAE itself results in average and blur generations and cannot reconstruct detailed textures of the masked region, and works poorly if we use MAE prior as the initial values for the stable diffusion model to perform image-to-image style inpainting, as illustrated in Figure 3. To this end, we adopt the MAE to provide prior to stabilizing SD.
Masking strategy for MAE. However, the original MAE is trained to estimate random masks uniformly distributed in the image. It is infeasible to directly adopt pre-trained MAE for inpainting task, which usually contains large continuous holes in the masked region. Inspired by Cao et al. [8], we fine-tune the MAE to inpainting masks. To adapt MAE for more practical inpainting scenarios, unlike Cao et al. [8] which only use continuous masks of range $[0.1,0.5]$, we construct a systematic masking strategy. Specifically, we generate masks with 50% object-shape mask, 20% Co-Mod mask [61], 20% LaMa [45] mask, 5% rectangle mask and 5% complement rectangle mask. We combine object-shape mask with Co-Mod style and LaMa style with the chance of 25%, respectively. This construction of mask style estimates the masks occurs in inpainting tasks, especially for the object removal and user-specified irregular masks. We control the mask ratio in the range of $[0.1,0.75]$ to follow the training scenario of MAE. For masks smaller than the ratio of 75%, we enlarge the mask ratio to 75% with randomly selected mask regions.
In practice, we train the MAE on images of size $256^{2}$, which is efficient and produce stable guidance for stable diffusion to generate high-resolution images.
3.2 Align MAE Prior with Stable Diffusion
Replace text-condition with MAE prior. SD is not trained on MAE priors. Fortunately, the text-condition part of SD is not utilized in inpainting, as we do not assume a text condition for inpainting task. Hence, we propose to replace the text-condition part of SD with our proposed MAE prior condition. However, as we do not fine-tune the SD, it cannot aligns well with the MAE prior. To achieve a proper guidance for SD, we introduces the alignment module to align both dimension and distribution.
Dimension alignment. Particularly, we adopt the MAE prior $F_{\mathrm{MAE}}$ of size $N_{m}\times M_{m}$, where $N_{m}$ is the sequence length and $M_{m}$ is the MAE feature dimension. To align it with the SD condition of size $N_{s}\times M_{s}$, we adopt a linear layer to map the feature dimension from $M_{m}$ to $M_{s}$ and set $N_{s}=N_{m}$ to preserve the local MAE prior.
Distribution alignment. After aligning the dimension, we use self-attention blocks to learn the distribution alignment for better guiding SD, leading to the condition $C_{\mathrm{MAE}}$ for SD. We train our alignment module with the same masking strategy used to train the MAE, while keeping the continuous-mask fine-tuned MAE and SD frozen. The training objective is the diffusion objective. To encourage the alignment module extract features from MAE prior, in the training stage, we use the full image to generate MAE prior instead of masked image with a chance of $p$.
Figure 4: The illustration of the information loss of KL-VAE. (a) The hue of the image is shifted. (b) KL-VAE still suffers from non-ignorable shifts in low-frequency fields.
3.3 Align Mask-Unmask Regions during Decoding
Information loss of KL-VAE. SD performs all its generative processes in the latent space and subsequently decodes these latent codes back to image space using KL-VAE. Despite the decoder being trained to reconstruct the image, it encounters challenges associated with information loss. Particularly in tasks like inpainting, the unmasked region necessitates replacement with ground-truth values. Though Rombach et al. [39] claimed that the diffusion model should prioritize the informative semantic compression, while the VAE and GAN are used to tackle perceptual compression with high-frequency details, we argue that low-frequency semantic loss in KL-VAE could not be neglected, as verified in Figure 4 (b). Moreover, as shown in Figure 4 (a), repeated decoding with KL-VAE not only noticeably degrades high-frequency details but also shifts in semantic hue. As human is sensitive to low-frequency information changes in the image, even subtle color shifts can induce significant perceptual inconsistencies. This results in low-fidelity inpainting results, especially under the irregular mask structures.
Figure 5: Training pipeline of the decoder, enhancing mask-unmask consistency via augmentations in color and latent spaces.
Figure 6: Qualitative SD1.5 inpainting results decoded by (b) vanilla KL-VAE decoder provided from [39], (c) larger decoder from [64], and (d) our augmented decoder. Our decoder can largely alleviate the mask-unmask color inconsistency.
Unmask-region conditioned decoder. The aforementioned information loss is prohibitive to be addressed by the SD U-Net alone, especially given the already degraded input latent codes. Inspired by Zhu et al. [64], we design a larger decoder to preserve the global consistency in low-frequency fields. This decoder, as shown in Figure 2, involves incorporating masked images in the pixel-wise color space as additional inputs. Unlike [64], we concatenate the 0-1 mask as the extra input to more precisely locate masked regions. However, the larger decoder with even pixel-wise image inputs still fails to handle the incompatible color and texture between the original images and compressed ones in challenging scenes as verified in Figure 6 (c). The low-frequency gap between degraded and original images makes it challenging to explicitly address this issue.
Mask-unmask consistent decoder. To train the decoder to ensure the consistency between generated latent and unmasked pixels, we estimate and enlarge the inconsistency during the training of the decoder. Specifically, we propose the color and latent augmentation as shown in Figure 5 to enhance the decoder to explicitly harmonize masked regions based on unmasked pixels.
i) Color augmentation: all training images would be augmented in brightness, contrast, saturation, and hue. This color augmentation is to avoid the decoder simply learns to lighter the image, as most VAE degradation results in a darker hue and oversaturation.
ii) Latent augmentation: To simulate the real VAE degradation, we propose leveraging the artifacts generated from the diffusion model. However, denoising to real images iteratively is notably time-consuming, even with DDIM [42]. Instead, we propose the one-step estimation of degraded latent $\tilde{\bm{z}}_{0}$ via the encoded clean latent $\bm{z}_{0}$ and noised latent $\bm{z}_{t}$:
$\tilde{\bm{z}}_{0}=\frac{1}{\sqrt{\overline{\alpha}_{t}}}(\hat{\bm{z}}_{t}-\sqrt{1-\overline{\alpha}_{t}}\varepsilon_{\theta}\left([\bm{z}_{t};\bm{z}_{0};\mathbf{M}],t\right)),$ (1)
where the timestep $t$ is randomly sampled from $[500,1000)$; $\overline{\alpha}_{t}$ indicates the cumulative production of prescribed variance schedule in diffusion; $\varepsilon_{\theta}(\cdot)$ is the noise estimation from frozen SD based on noised $\bm{z}_{t}$, unmasked $\bm{z}_{0}$, and all-zero masking $\mathbf{M}$. Despite $\tilde{\bm{z}}_{0}$ suffering from an extremely large step denoising, we empirically find that the inpainting fine-tuned SD could produce reliable results with little artifacts enhanced by the concatenation of unmasked latent $\bm{z}_{0}$. Moreover, the latent augmentation makes the image undergo two-round encoding and decoding from KL-VAE. This not only facilitates the incorporation of natural VAE degradation into the decoder training but also enhances the overall robustness of the model. We empirically apply 50% to operate the latent augmentation to training images. We fine-tune the decoder based on [64], showcasing superior consistency as compared in Figure 6.
4 Evaluation Dataset
Figure 7: Examples of masked images constructed in MISATO. The first three cases focus on different image domains, while in the last case we focus on the inpainting of background.
To validate the inpainting performance across different domains and mask styles, we construct a evaluation dataset, dubbed as MISATO, from Matterport3D [10], Flickr-Landscape [29], MegaDept[27], and COCO 2014 [30]. Matterport3D contains a indoor image dataset, Flickr-Landscape is consisted of outdoor landscape images, MegaDepth involves a dataset of buildings, and COCO has precise segmentation of foreground subjects. For each dataset, we select 500 representative examples, yielding 2000 testing examples of size $512^{2}$.
The principle of constructing MISATO is to select the most representative and diverse examples. To this end, for Matterport3D, Flickr-Landscape and MegaDepth, we use CLIP visual model [37] to extract semantic visual features. Then we use BisectingKMeans [44] to cluster each dataset into 500 clusters, and select the cluster centers as the evaluation data. The selected data are center cropped and then resized to $512^{2}$. For COCO, we focus on the background inpainting. To this end, for each data we identify the foreground with provided segmentation and remove it from the generated masks, yielding a dataset specified for purely background inpainting.
Combined together, MISATO contains four inpainting domains, indoor, outdoor landscape, building, and background. As shown in Figure 7, for each evaluation image, we adopt the similar masking strategy as in Section 3.1 while not using the rectangle and complement rectangle masks. The masking ratio is set as $[0.2,0.8]$.
5 Experiments
Implementation details. ASUKA utilizes the stable diffusion text-guided inpainting model v1.5 as the generative backbone. We train ASUKA on Places2 [63]. For the MAE [20] used in ASUKA, we fine-tune from the MAE used in Cao et al. [8] with a batch size of 1024. We use four self-attention blocks for the alignment module, training from scratch with AdamW [32] optimizer of learning rate 5e-2 for 23K steps of batch size 64 with the standard diffusion objective. We use warm-up strategy to set initial $p$ as $100\%$ and linearly decay it to $10\%$ in the first 2k training steps and then freeze $p=10\%$. We fine-tune the decoder based on [64] for 50K steps with a batch size of 40 and learning rate of 8e-5 with cosine decay. We use ColorJitter to implement color augmentation, with the setting of brightness 0.15, contrast 0.2, saturation 0.1, and hue 0.03.
Evaluation metrics. We compare inpainting algorithms using the following metrics: Peak Signal-to-Noise Ratio (PSNR) to compare the ratio of real data to introduced noise, Structural Similarity Index Measure (SSIM) [47] to identify the perceived change in structural information, Learned Perceptual Image Patch Similarity (LPIPS) [60] to calculate the patch-level image distances, Fréchet Inception Distance (FID) [21] to compare the distribution distance between generated images and real images, and Paired/Unpaired Inception Discriminative Score (P-IDS/U-IDS) [61] to measure the human-inspired linear separability.
5.1 Comparison on MISATO
Figure 8: Qualitative comparison on MISATO. We compare ASUKA with Co-Mod [61], MAT [26], LaMa [45], MAE-FAR [8], and SD [39] in landscape (first two rows), indoor (3rd and 4th rows), building (5th and 6th rows) and background inpaintings (last two rows). SD performs unconditional generation. SD-text utilizes text prompt of “background”. SD-prompt utilizes trained prompt for inpainting task using the same training setting of ASUKA. Inpainting algorithms usually suffer from fidelity issue, while SD-based algorithms suffer from stability issue and generate random elements. In contrast, ASUKA achieves stable and faithful inpainting.
[…]
Co-Mod [61] 21.805 0.787 0.179 17.421 0.243 0.109
MAT [26] 21.553 0.787 0.176 17.261 0.255 0.122
LaMa [45] 23.677 0.820 0.155 15.436 0.260 0.135
MAE-FAR [8] 24.150 0.821 0.142 13.283 0.282 0.153
SD [39] 22.454 0.796 0.168 12.812 0.345 0.211
SD-text 22.394 0.793 0.164 12.603 0.337 0.207
SD-prompt 22.381 0.798 0.160 12.517 0.331 0.204
ASUKA 22.495 0.799 0.150 11.460 0.368 0.256
Table 1: Quantitative comparison on MISATO. Co-Mod [61], MAT [26], LaMa [45], and MAE-FAR [8] are state-of-the-art inpainting algorithms. SD [39] performs unconditional generation. SD-text utilizes text prompts of “background” to perform text-guided generation. SD-prompt utilizes learnable prompts trained specifically for inpainting, using the same training setting as ASUKA, performing prompt-guided generation. ASUKA and SD use the stable diffusion text-guided inpainting model v1.5.
Competitors. We first compare ASUKA with SD [39]. Specifically, we consider three variants of SD: i) SD, which generates inpainting results with no guidance, and performs unconditional generation; ii) SD-text, which generates inpainting results guided by the text prompt of “background”; iii) SD-prompt, where we replace the text prompt with learnable tokens and train the tokens using the same training pipeline of ASUKA. Beyond SD, we also compare ASUKA with state-of-the-art inpainting algorithms, including Co-Mod [61], MAT [26], LaMa [45], and MAE-FAR [8].
Quantitative comparison results are reported in Table 1. ASUKA, although based on frozen SD, achieves significant and consistent improvements over SD across all the evaluation metrics, and enjoys the state-of-the-art results on FID, U-IDS and P-IDS. Particularly, the U-IDS and P-IDS coincide with human preference [61], and has a potential upper bound of $0.5$, indicating the superiority of ASUKA by a large margin. The inconsistency results between PSNR, SSIM, LPIPS and FID, U-IDS, P-IDS, partially comes from the frozen SD, where ASUKA achieves consistent improvements but is bounded to the frozen U-Net. On the other hand, the PSNR and SSIM is known to have bad correlation with human preferences, as several illustrative examples shown in the appendix. In summary, the results validate the contribution of ASUKA that yields stable and faithful inpainting through frozen SD.
Qualitative comparison examples are shown in Figure 8. i) The state-of-the-art inpainting algorithms usually suffer from fidelity issues, for example the unnatural boundaries in the third and fourth rows, and failed inpainting of tower in the third-to-last row. LaMa and MAE-FAR sometimes lead to blurred inpainting results, especially in the scenario of large continuous masks. ii) The three variants of SD usually suffer from the stability issue and generate random new elements in the masked region, in almost all the illustrated images. iii) Particularly, for the background mask shown in the second-to-last row, state-of-the-art inpainting algorithms cannot identify the object precisely, and extend the human in hairs and elbows. On the other hand, SD generates random elements in the masked region. iv) In contrast, ASUKA enjoys superior stable and faithful inpainting.
5.2 Further Analysis of ASUKA
[…]
p.t. MAE 22.485 0.799 0.151 11.513 0.354 0.258
f.t. MAE 22.495 0.799 0.150 11.460 0.368 0.256
Table 2: Quantitative comparison of ASUKA using pre-trained MAE (p.t. MAE) from [8] and fine-tuned MAE (f.t. MAE).
Ablation of MAE prior. We compare our fine-tuned MAE with directly adopting the MAE trained in [8] without further fine-tuning. To this end, we train ASUKA with the MAE in [8] using the same training strategy and compare the results in. Results suggest the improvements of fine-tuning MAE, especially on FID and U-IDS. This improvement comes from the better adaptation on the real-world masks.
[…]
linear 22.402 0.797 0.155 11.934 0.361 0.227
attn 22.474 0.798 0.152 11.613 0.362 0.234
cross x4 22.488 0.799 0.152 11.762 0.368 0.238
ASUKA 22.495 0.799 0.150 11.460 0.368 0.256
Table 3: Ablation of different alignment modules. Linear adopts linear layer to align feature dimension only; attn adopts a self-attention layer to further align distribution; cross x4 adopts 4 cross-attention layers; while ASUKA adopts 4 self-attention layers.
Ablation of alignment module. Our alignment module includes a dimension and a distribution alignment. We validate the effectiveness of each module by comparing with the following variants step by step: i) linear: Use linear layer to align feature dimension only; ii) attn: Based on i, further use a single self-attention block to align the distribution; iii) cross x4: we instead use learnable query and 4 cross-attention layers to learn the MAE prior. ASUKA adopts 4 self-attention blocks. Results are shown in Table 3. The self-attention block shows improved results on simply aligning the dimension as well as using a cross-attention block to align the distribution. Using 4 self-attention blocks improves the capacity of the alignment module.
[…]
KL-VAE [39] 22.454 0.795 0.156 11.949 0.343 0.208
Asy. VAE [64] 22.566 0.801 0.151 11.634 0.361 0.231
color-aug 22.475 0.799 0.152 11.603 0.357 0.229
Ours 22.495 0.799 0.150 11.460 0.368 0.256
Table 4: Quantitative comparison of different decoders for SD.
Ablation of decoder training pipeline. For the decoder, we compare ASUKA with i) KL-VAE: the decoder used in SD; ii) Asy. VAE: the decoder trained by [64]; iii) color-aug: only use color augmentation to train the decoder; Results are shown in Table 4, which shows the superiority of our augmented decoder to ensure the mask-unmask consistency.
Figure 9: The curse of self-attention, causing the MAE falsely estimate the masked region and powerful text-guided diffusion models fail to generation content based on text prompts. ASUKA potential circumvents this issue by using a blank paper image as the input to the MAE to provide correct prior.
Limitation: The “curse” of self-attention. The primary limitation of ASUKA stems from the inefficacy of the MAE prior, mainly due to issues within the self-attention module. Specifically, as showin in Figure 9, the presence of multiple similar objects in an image may lead the MAE to incorrectly predict a similar object in the masked region, conflicting with the goal of object removal. Notably, this curse of self-attention significantly impacts diffusion-based generative models. It results in the inability to accurately follow “blank paper” text prompts, even when employing a substantial classifier-free guidance scale of 9. This issue is not unique to SD but is also a common problem in other advanced text-guided diffusion models, such as OpenAI’s DALL-E 2 [38] and Adobe’s FileFly [1]. Nevertheless, ASUKA has the potential to circumvent this issue by modifying the MAE prior, for instance, by instead using a blank paper image as the input to MAE prior. A more comprehensive solution would involve extra control on self-attention layers in diffusion models, which we leave as future work.
6 Conclusion
In this paper, we proposed Aligned Stable inpainting with Unknown Areas prior (ASUKA) to achieve stable and faithful inpainting via frozen Stable Diffusion inpainting model (SD). To achieve stability, we adopt a reconstruction-based masked-auto encoder (MAE) as the stable prior for masked region purely from unmasked region. Then we align the stable prior to frozen SD with the proposed alignment module. To achieve fidelity, we resolve the mask-unmask inconsistency of SD in the decoding process. We fine-tune a unmask-region conditioned KL-VAE decoder to perform local harmonization during the decoding process. To validate the efficacy of inpainting algorithms in different image domains and mask types, we introduce an evaluation dataset, named as MISATO, from existing datasets. ASUKA enjoys superior stable and faithful inpainting results than state-of-the-art inpainting algorithms and SD-based variants.
Supplementary Material
7 Extra Analysis
7.1 Comparison on Places2
[…]
Co-Mod [61] 19.413 0.709 0.267 5.794 0.274 0.096
MAT [26] 19.729 0.740 0.202 3.765 0.348 0.195
LaMa [45] 21.503 0.771 0.202 6.693 0.247 0.050
MAE-FAR [8] 22.140 0.779 0.174 3.559 0.307 0.105
SD [39] 20.302 0.746 0.193 1.514 0.375 0.207
SD-text 20.359 0.748 0.191 1.506 0.373 0.202
SD-prompt 20.548 0.753 0.189 1.477 0.390 0.234
ASUKA 20.406 0.752 0.183 1.230 0.413 0.287
Table 5: Quantitative comparison on Places 2.
We evaluation ASUKA and other inpainting algorithms on the standard testing benchmark dataset of Places2 [63], which includes 36500 testing examples. Results on this dataset also validate the efficacy of ASUKA.
7.2 User-Study
Model Consistency (%) Quality (%)
Co-Mod [61] 4.69 5
MAT [26] 7.5 3.33
LaMa [45] 8.44 8.33
MAE-FAR [8] 5.31 5.83
SD [39] 10.63 5.83
SD-text 7.81 15.83
SD-prompt 16.25 15.83
ASUKA 39.38 40
Table 6: User-study of top-1 ratio on consistency and quality among all the inpainting results.
To evaluate the user preference on inpainting algorithms, we conduct an user-study. Specifically, we randomly select 40 testing images along with inpainting results on these images. We ask the user to select the best inpainting results from the following perspectives respectively: i) Consistency: the masked region should be consistent with unmasked region, involving color and texture consistency and not generating new elements; ii) Quality: the fidelity of the inpainting results. The first metric measures the stability while the other measures the fidelity of the inpainting results. We collect 100 valid anonymous questionnaire results, and report the average selection ratio among all the inpainting algorithms. This result validate the efficacy of ASUKA on alignment with human preference.
7.3 The Misalignment Between PSNR, SSIM and Human Preference
The PSNR and SSIM are known to be have bad correlation with human preferences. To show this, we report the PSNR and SSIM scores for the qualitative comparison on MISATO ( Figure 8 in the paper) in Figure 10. Higher PSNR and SSIM do not always indicate better results that human prefers. Particularly, blurred inpainting results usually have higher PSNR and SSIM scores, like the 1st, 2rd, 4th, 8th rows. However, it is clear that blurred results are not expected.
Figure 10: Misalignment between PSNR, SSIM and human preference. Images are same with Figure 8. PSNR and SSIM scores are listed in the bottom, where the best values are marked as red.